OpenStudy (anonymous):

given that y varies directly with x write the linear function relating the two variables if y=2 and x=10

OpenStudy (saifoo.khan):

y = kx Where k is the constant.

OpenStudy (saifoo.khan):

Now find the value of the constant using the values given, y = kx 2 = k(10) 2/10 = k k = 1/5

OpenStudy (anonymous):

see i Read hat but it is not clicking for me as what to do next to solve the dang problem

OpenStudy (anonymous):

oh, I got too hasty sorry

OpenStudy (saifoo.khan):

No worries! ^_^

OpenStudy (anonymous):

thank you Saifoo.kahn

OpenStudy (saifoo.khan):

No problem! Have more? Feel free to post them! :)